Davis Industries must choose between a gas-powered and an electric-powered forklift truck for moving materials in its factory. Because both forklifts perform the same function, the firm will choose only one. (They are mutually exclusive investments.) The electric-powered truck will cost more, but it will be less expensive to operate; it will cost $21,000, whereas the gas-powered truck will cost $17,230. The cost of capital that applies to both investments is 11%. The life for both types of truck is estimated to be 6 years, during which time the net cash flows for the electric-powered truck will be $6,100 per year and those for the gas-powered truck will be $5,300 per year. Annual net cash flows include depreciation expenses.

  1. Calculate the NPV for each type of truck. Do not round intermediate calculations. Round your answers to the nearest dollar.

    Electric-powered forklift truck: ____ & Gas-powered forklift truck: ______

  2. Calculate the IRR for each type of truck. Do not round intermediate calculations. Round your answers to two decimal places.

    Electric-powered forklift truck: ____ & Gas-powered forklift truck: ______
Answer #1

a. NPV of Electric-powered forklift truck= PV of net cash flows -Initial Investment =6100*((1-(1+11%)^-6)/11%-21000=4806.28
NPV of Gas-powered forklift truck=PV of net cash flows -Initial Investment =5300*((1-(1+11%)^-6)/11%-17230=5191.85

b. NPV of Electric-powered forklift truck
Using Financial calculator to calculate IRR
N=6;PMT=6100;PV=-21000;CPT IRR =18.62%
NPV of Gas-powered forklift truck
N=6;PMT=5300;PV=-17230;CPT IRR =20.92%